Ql: For the frame shown in the Fig. (1), check the determinacy and stability then draw the axial force, shearing force, and bending moment diagrams for the frame.

1.2 k/ft Pin 25 k 15 ft 15 k B 20 ft 15 ft А 30 ft — Fig. (1) Fi

I need it in 30min or 1 h
Answer #1

We need to check the determinacy of frame so by using formula Ds= 3m + r -( 3j + no. of independent equation)

here m is no. of members.=3

. r is no. of reaction.=4,

j is no. of joints = 4

Ds= 3*3+4 -(3*4 + 1) = 0 so the structure is determinate.

Rest question is solved in the following images.

keeping in cao HATHA-10 Ha=-10-15 --- 7.5k. S ve sign shows reverse dis HA=-17.5k Now from left side. MC=o. HAX 30 + VAX 30 2BMD MB = 350 Kft to cation of length of B BC - 316aft 8.83 27017 31.62.2 7.766 Q = 7 16 full Horizontal distance ga. Cos 18°